OpenStudy (across):

Here is a fun exercise: \[\log_{2}7.\]Prove that the above expression is irrational.

hero (hero):

Why is it always that a proof is considered a "fun" exercise?

OpenStudy (anonymous):

to try to get someone to try it :)

OpenStudy (anonymous):

7 <> 2^p/2^q where p and q are integers ?

OpenStudy (across):

It's not immediately obvious 2^p/2^q ≠ 7, but you're on the right track. Perhaps I should post the proof?

hero (hero):

Yeah, that would be helpful

OpenStudy (across):

Proof by contradiction: Assume that the expression is rational. This implies that \[\log_{2}7=\frac{p}{q},\]\[p, q \in \mathbb{Z},\]\[q \neq 0.\]By simple arithmetic, \[\log_{2}7=\frac{p}{q},\]\[q\log_{2}7=p,\]\[\log_{2}7^{q}=p,\]\[7^{q}=2^{p}.\]Note that the LHS of this expression is always odd whilst its RHS is always even. This is a contradiction since an odd number cannot be equal to an even number, and therefore, \[\log_{2}7\]is irrational. ∎

OpenStudy (anonymous):

nice, thanks :)

OpenStudy (anonymous):

really good

OpenStudy (anonymous):

classic proof by contradiction

hero (hero):

I hate proof by contradiction